OpenStudy (anonymous):

The jet left paris at the same time as a pasenger plane. the plane flew in the opposite directions. The pasenger flew at a speed of 450mph. After 11 hours they were in 9460mi. How fast did the jet fly?

OpenStudy (anonymous):

D of passenger plane after 11 hours is \[450*11=4950\]D of jet is \[11x\]Total distance = 4950 + 11x = 9460 x = 410